OpenStudy (anonymous):

-5y+3y=0 Find the slope and y-intercept

OpenStudy (anonymous):

slope of \(ax+by=c\) is \(-\frac{a}{b}\) in your case it is \(-\frac{-5}{3}=\frac{5}{3}\) assuming you meant \[-5x+3y=0\]

OpenStudy (anonymous):

i imagine \(-5y+3y=0\) is a typo